# On the cardinality of measures of maximal relative entropy for smooth skew products

### v2

@@ 1.3. An application of Corollary B
$T(x_{1},x_{2})=(2x_{1}-x_{2}+k\cos(2\pi x_{1}),\,x_{1})\pmod{1},\ \text{for some }k\in\mathbb{Z}.$
Sinai conjectured that for $k$ sufficiently large, $T$ has positive metric entropy with respect to Lebesgue measure [47, Page 144]. This conjecture remains open. In contrast, much more is known about measures of maximal entropy: Obata showed that, for large $k$, the standard map admits a unique measure of maximal entropy; moreover, such a measure has positive metric entropy [40]. However, less is known in the random setting. Our results show that there exists at least one and, at most, countably many ergodic relative measures of maximal entropy for certain classes of random perturbations of the standard map.
For each $k\in\mathbb{N}$, we consider the skew product
@@
Let $\mathbb{P}$ be an ergodic $\theta$-invariant measure. Then, there exists $k_{0}\in\mathbb{N}$, such that for every $k\geq k_{0},$ one has $h_{\mathrm{top}}(\Theta_{k}\mid\mathbb{P})>0$. In particular, Proposition 1.5 implies that $\Theta_{k}$ admits a measure of maximal $\mathbb{P}$-relative entropy and Corollary B ensures that there are only countably many such ergodic measures.
@@ 2. Hyperbolic measures and Borel homoclinic classes
In this section, we collect standard facts from hyperbolic dynamics and introduce the concept of Borel homoclinic classes (see [19, Definition 2.7]). Fix a compact smooth manifold $E$ and a $\mathcal{C}^{1+\alpha}$ diffeomorphism $\Theta:E\to E$, for some $\alpha>0$. Unless stated otherwise, all statements in this section are made under this standing assumption. For the purposes of this paper, the set $E$ can be thought of as $\Omega\times M$ as in Theorem A. Our exposition closely follows [19, Section 2].
@@ 2.1. Hyperbolic measures
By the Oseledets’ multiplicative ergodic theorem [41], there exists a Borel set $E^{\prime}\subset E$ such that
@@
    $\sigma^{+}(x):=\{\lambda\in\sigma(x):\lambda>0\}\neq\ \emptyset\quad\text{and}\ \sigma^{-}(x):=\{\lambda\in\sigma(x):\lambda<0\}\neq\emptyset.$
@@
Finally, we recall that if (i) holds but (ii) fails, then $\mu$ is supported on a periodic orbit and hence has zero metric entropy [27, Corollary S.5.2] (see also the proof of [26, Theorem 4.2]).
@@
Given $\chi,\varepsilon>0$, a $(\chi,\varepsilon)$-Pesin block is a non-empty set $\Lambda\subset E$ for which there exists a direct sum $T_{x}E=E^{s}(x)\oplus E^{u}(x)$ for all $x\in\bigcup_{n\in\mathbb{Z}}\Theta^{n}(\Lambda),$ and a number $K>0$ such that for any $n\in\mathbb{Z}$, $k\geq 0,$ and $y\in\Lambda$,
@@
Below we state Pesin’s stable manifold theorem, whose proof can be found at [43] (see also [19, Theorem 2.2]).
@@
We recall that the choice of $\varepsilon>0$ in Theorem 2.3 does not affect the definition of the global stable and unstable manifolds [9, §8.2] (see also [19, Section 2]).
@@ 2.2. Borel homoclinic classes
Definition 2.6 ([19], Homoclinic relation).
@@
Proposition 2.7 ([19, Proposition 2.6]).
@@
Proposition 2.9 ([19, Proposition 2.8]).
@@
Note that any Borel homoclinic class that is not a singleton supports uncountably many ergodic invariant measures. If $\Theta$ admits a non-atomic hyperbolic invariant measure (i.e., not supported on a periodic orbit), then $\Theta$ admits uncountably many hyperbolic ergodic measures. Indeed, every non-trivial Borel homoclinic class contains at least one horseshoe, and each horseshoe carries uncountably many ergodic measures since it is conjugated to a full shift. By Proposition 2.9 (iv), each of these measures is hyperbolic.
@@ 3. Topological Markov shifts
equipped with the metric $d(v,u)=\min\{\exp(-\min\{|n|;v_{n}\neq u_{n}\})\}$, and $\sigma:\Sigma\to\Sigma$ is the left shift map $\sigma((v_{i})_{i\in\mathbb{Z}})=(v_{i+1})_{i+1}$. Given a topological Markov shift $\Sigma$ associated to a graph $\mathscr{G}=(\mathscr{V},\mathscr{E})$ we define the set
@@
Definition 3.2 (Bowen Partition).
Let $\theta:\Omega\to\Omega$ be a homeomorphism. We say that $\theta$ admits a Bowen partition if there exist a family of compact sets $\mathscr{P}\subset\mathcal{B}(\Omega)$ such that
@@
    $\mathscr{P}<\infty$;
@@
    let $\mathscr{G}_{b}=(\mathscr{V}_{b},\mathscr{E}_{b})$, with $\mathscr{V}_{b}=\mathscr{P}$ and
    $\mathscr{E}_{b}=\{P\to P^{\prime};\,P,P^{\prime}\in\mathscr{P}\ \text{and }P\cap\theta^{-1}(P^{\prime})\neq\emptyset\}.$
    Setting the subshift of finite type $\Sigma_{b}=\Sigma_{b}(\mathscr{G}_{b})$ the have that map
    $\pi_{b}:(P_{i})_{i\in\mathbb{Z}}\in\Sigma_{b}\mapsto\bigcap_{i\in\mathbb{Z}}\theta^{-i}(P_{i})\in\Omega,$ (3)
    is well-defined and Hölder continuous; and
@@
    the function $\pi$ is surjective over the non-wandering set $NW(\theta)$ and $\pi^{-1}(\omega)\leq k_{\mathscr{P}}$ for some $k_{\mathscr{P}}\in\mathbb{N}$ for every $\omega\in NW(\theta)$.
Observe that under the above assumptions imply that $\pi_{b}\circ\sigma=\theta\circ\pi_{b}.$
Theorem 3.3 ([14], [42, Appendix III]).
Every $\mathcal{C}^{1+\alpha}$Axiom $A$ diffeomorphism admits a Bowen partition.
The next statement, taken from [19, Theorem 10.5 and Section 10], is a higher-dimensional generalisation of Sarig’s celebrated theorem [45, Theorem 1.3] (see also [10, Theorem 0.1]). We mention that the original result presented in [19] is more general; however, the presented version is enough for our purposes.
Theorem 3.4 ([19, Theorem 10.5 and Section 10]).
Let $\Theta:E\to E$ be a $\mathcal{C}^{1+\alpha}$ diffeomorphism on a compact smooth manifold $E$. Let $\mathcal{X}$ be either a Borel homoclinic class $H$ or $E$. Then, for each $\chi>0$, there exists a countable collection of sets $\mathscr{R}\subset\mathscr{B}(M)$, such that the topological Markov shift $\Sigma=\Sigma(\mathscr{G})$, induced by the graph $\mathscr{G}=(\mathscr{V},\mathscr{E})$ with $\mathscr{V}:=\mathscr{R}$ and $\mathscr{E}=\{R\to R^{\prime};R,R^{\prime}\in\mathscr{R}\ \text{and }R\cap\Theta^{-1}(R^{\prime})\neq\emptyset\},$ satisfies the following properties:
@@
    $\pi:(R_{i})_{i\in\mathbb{N}}\in\Sigma\mapsto\bigcap_{n\in\mathbb{N}}\overline{\bigcap_{-n\leq i\leq n}\Theta^{-i}(R_{i})}\in\mathcal{X}$
    is well defined and Hölder continuous. Moreover, $\Theta\circ\pi=\pi\circ\sigma$ and $\pi:\Sigma^{\#}\to\mathcal{X}$ is finite-to-one (see Definition 3.1).
@@
    $\Sigma$ is a locally compact countable-state Markov shift and $\mu(\pi(\Sigma))=\mu(\pi(\Sigma^{\#}))=1$. If $\mathcal{X}$ is a Borel homoclinic class, we can choose $\Sigma$ as irreducible;
@@
    Every ergodic invariant probability measure $\widehat{\mu}$ on $\Sigma$ projects to an ergodic $\Theta$-invariant measure $\mu:=\pi_{*}\widehat{\mu}$ such that $\mu(\mathcal{X})=1$ and $h_{\mu}(\Theta)=h_{\widehat{\mu}}(\sigma).$
@@
    For every $\chi$-hyperbolic $\Theta$-invariant probability measure $\mu$ on $M$ there exists a $\sigma$-invariant probability measure $\widehat{\mu}$ on $\Sigma$ such that $\pi_{*}\widehat{\mu}=\mu$.
We call $\mathscr{R}$ a $\chi$-Sarig partition of $\Theta$.
We will prove the following theorem:
Our results are based on the following theorem, which is based on a refinement of Theorem 3.4 for skew products.
@@
    $\Theta$ is a skew product i.e. $\Theta(\omega,x)=(\theta(\omega),T_{\omega}(x))$; and
@@
    $\theta:\Omega\to\Omega$ is Axiom $A$.
Let $\mathcal{X}\subset\Omega\times M$ be equal to a Borel homoclinic class $H$ or $\Omega\times M$. Then, for every $\chi>0$ there exist a locally compact topological Markov shift $\Sigma_{s}$, a subshift of finite type $\Sigma_{b}$ and Hölder continuous maps $\pi_{s}:\Sigma_{s}\to\mathcal{X}$ and $\pi_{b}:\Sigma_{b}\to\Omega$ such that
@@
    for every $\chi$-hyperbolic $\Theta$-invariant probability $\mu$ such that $\mu(\mathcal{X})=1$, one has $\mu\big(\pi_{s}[\Sigma_{s}]\big)=\mu\big(\pi_{s}[\Sigma_{s}^{\#}]\big)=1$};
@@
    the maps $\pi_{s}|_{\Sigma_{s}^{\#}}:\Sigma_{s}^{\#}\to\mathcal{X}$ and $\pi_{b}:\Sigma_{b}\to\Omega$ are finite-to-one;
@@
    there exists a one $1$-block factor map $\mathrm{proj}_{\Sigma_{b}}:\Sigma_{s}\to\Sigma_{b}$ such that
@@
    Write $\Sigma_{b}^{\prime}:=\mathrm{proj}_{\Sigma_{b}}(\Sigma_{s})$. If $\mathcal{X}=\Omega\times M$ then $\mathrm{proj}_{\Sigma_{b}}$ is onto and $\Sigma_{b}^{\prime}=\Sigma_{b}$. If $\mathcal{X}$ is a Borel homoclinic class, then $\Sigma_{b}^{\prime}$ is a subshift of finite type; and
@@
    if $\mathcal{X}$ is a Borel homoclinic class, $\Sigma_{s}$ can be chosen irreducible; in this case $\Sigma_{b}^{\prime}$ is an irreducible subshift of finite type.
@@
$\mathcal{X}\subset\Omega\times M$$\mathcal{X}\subset\Omega\times M$$\Sigma$$\Sigma$$\Sigma_{s}$$\Sigma_{s}$$\Sigma_{b}^{\prime}\subset\Sigma_{b}$$\Sigma_{b}^{\prime}\subset\Sigma_{b}$$NW(\theta)\subset\Omega$$NW(\theta)\subset\Omega$$\scriptstyle{{\rm proj}_{\Omega}}$$\scriptstyle{\Theta}$$\scriptstyle{{\rm proj}_{\Omega}}$$\scriptstyle{\sigma}$$\scriptstyle{\pi}$$\scriptstyle{\pi}$$\scriptstyle{\pi_{s}}$$\scriptstyle{{\rm proj}_{\Sigma_{b}}}$$\scriptstyle{\sigma_{s}=\sigma_{b}\times\sigma}$$\scriptstyle{{\rm proj}_{\Sigma}}$$\scriptstyle{{\rm proj}_{\Sigma}}$$\scriptstyle{\pi_{s}}$$\scriptstyle{{\rm proj}_{\Sigma_{b}}}$$\scriptstyle{\sigma_{b}}$$\scriptstyle{\pi_{b}}$$\scriptstyle{\pi_{b}}$$\scriptstyle{\theta}$
@@
Let $\chi>0$. Applying Theorem 3.4 to the skew product $\Theta:\Omega\times M\to\Omega\times M$ and $\mathcal{X}$ be either equal to Borel homoclinic class $H$ or $\Omega\times M$, we obtain a topological Markov shift $(\Sigma,\sigma)$ with the properties asserted therein, and let $\mathscr{R}$ be its $\chi$-Sarig partition. We assume $(\Sigma,\sigma)$ to be irreducible if $\mathcal{X}$ is a Borel homoclinic class. Since $\theta:\Omega\to\Omega$ is an Axiom A diffeomorphism, there exists a subshift of finite type $(\Sigma_{b},\sigma_{b})$ associated with a Bowen partition $\mathscr{P}$ as in Definition 3.2, let $\pi_{b}$ be as in (3) and recall that
$k_{\mathscr{P}}:=\sup_{\omega\in NW(\theta)}\#\pi_{b}^{-1}(\omega)<\infty.$
Define the graph $\mathscr{G}_{s}:=(\mathscr{V}_{s},\mathscr{E}_{s})$ where
$\mathscr{V}_{s}:=\{(P,R);\,R\in\mathscr{R},P\in\mathscr{P}\ \text{and }\overline{R}\cap(P\times M)\neq\emptyset\},$

### v3

@@ 1.3. An application of Corollary B
$T(x_{1},x_{2})=(2x_{1}-x_{2}+k\cos(2\pi x_{1}),\,x_{1})\pmod{1},\ \text{for some }k>0.$
Sinai conjectured that for $k$ sufficiently large, $T$ has positive metric entropy with respect to Lebesgue measure [Sinai1994, Page 144]. This conjecture remains open. In contrast, much more is known about measures of maximal entropy: Obata showed that, for large $k$, the standard map admits a unique measure of maximal entropy; moreover, such a measure has positive metric entropy [Obata2021]. However, less is known in the random setting. Our results show that there exists at least one and, at most, countably many ergodic relative measures of maximal entropy for certain classes of random perturbations of the standard map.
For each $k>0$, we consider the skew product
@@
Let $\mathbb{P}$ be a $\theta$-invariant ergodic measure with full support. Then, there exists $k_{0}>0$, such that for every $k>k_{0},$ one has $h_{\mathrm{top}}(\Theta_{k}\mid\mathbb{P})>0$. In particular, Proposition 1.5 implies that $\Theta_{k}$ admits a measure of maximal $\mathbb{P}$-relative entropy and Corollary B ensures that there are only countably many such ergodic measures.
@@ 2. Hyperbolic measures and Borel homoclinic classes
In this section, we collect standard facts from hyperbolic dynamics and introduce the concept of Borel homoclinic classes (see [buzzi2025strongpositiverecurrenceexponential, Definition 2.7]). Fix a compact smooth manifold $E$ and a $\mathcal{C}^{1+\alpha}$ diffeomorphism $\Theta:E\to E$, for some $\alpha>0$. Unless stated otherwise, all statements in this section are made under this standing assumption. For the purposes of this paper, the set $E$ can be thought of as $\Omega\times M$ as in Theorem A. Our exposition closely follows [buzzi2025strongpositiverecurrenceexponential, Section 2].
@@ 2.1. Hyperbolic measures
By the Oseledets’ multiplicative ergodic theorem [Oseledets68], there exists a Borel set $E^{\prime}\subset E$ such that
@@
    $\sigma^{+}(x):=\{\lambda\in\sigma(x):\lambda>0\}\neq\ \varnothing\quad\text{and}\ \sigma^{-}(x):=\{\lambda\in\sigma(x):\lambda<0\}\neq\varnothing.$
@@
Finally, we recall that if (i) holds but (ii) fails, then $\mu$ is supported on a periodic orbit and hence has zero metric entropy [KatokBook, Corollary S.5.2] (see also the proof of [Katok80, Theorem 4.2]).
@@
Given $\chi,\varepsilon>0$, a $(\chi,\varepsilon)$-Pesin block is a nonempty set $\Lambda\subset E$ for which there exists a direct sum $T_{x}E=E^{s}(x)\oplus E^{u}(x)$ for all $x\in\bigcup_{n\in\mathbb{Z}}\Theta^{n}(\Lambda),$ and a number $K>0$ such that for any $n\in\mathbb{Z}$, $k\geq 0,$ and $y\in\Lambda$,
@@
Below we state Pesin’s stable manifold theorem, whose proof can be found at [Pesin] (see also [buzzi2025strongpositiverecurrenceexponential, Theorem 2.2]).
@@
We recall that the choice of $\varepsilon>0$ in Theorem 2.3 does not affect the definition of the global stable and unstable manifolds [PesinBook, §8.2] (see also [buzzi2025strongpositiverecurrenceexponential, Section 2]).
@@ 2.2. Borel homoclinic classes
Definition 2.6 ([buzzi2025strongpositiverecurrenceexponential], Homoclinic relation).
@@
Proposition 2.7 ([buzzi2025strongpositiverecurrenceexponential, Proposition 2.6]).
@@
Proposition 2.9 ([buzzi2025strongpositiverecurrenceexponential, Proposition 2.8]).
@@
Note that any Borel homoclinic class that is not a singleton supports uncountably many invariant ergodic measures. If $\Theta$ admits a non-atomic hyperbolic invariant measure (i.e., not supported on a periodic orbit), then $\Theta$ admits uncountably many hyperbolic ergodic measures. Indeed, every non-trivial Borel homoclinic class contains at least one horseshoe, and each horseshoe carries uncountably many ergodic measures since it is conjugated to a full shift. By Proposition 2.9 (iv), each of these measures is hyperbolic.
We conclude this section by introducing the notions of rectangle with local product structure and the Smale bracket, which will be used throughout.
Definition 2.11 (Rectangle and Smale bracket).
Let $\Theta:E\to E$ be a dynamical system on a smooth manifold $E$. A set $R\subset E$ is a rectangle with local product structure if there exist families $\Gamma^{s}$, $\Gamma^{u}$ of $\mathcal{C}^{1}$ discs embedded in $E$ such that:
  • $R=\bigcup_{\gamma^{s}\in\Gamma^{s}}\gamma^{s}\cap\bigcup_{\gamma^{u}\in\Gamma^{u}}\gamma^{u}$;
  • given $\gamma,\gamma^{\prime}\in\Gamma^{s}$ (or $\Gamma^{u}$), then either $\gamma=\gamma^{\prime}$ or $\gamma\cap\gamma^{\prime}=\emptyset.$
  • each $\gamma^{s}\in\Gamma^{s}$ is a piece of stable manifold and for each $x\in R$, there exists a unique $\gamma^{s}_{x}\in\Gamma^{s}$ such that $x\in\gamma^{s}_{x}$;
  • each $\gamma^{u}\in\Gamma^{u}$ is a piece of unstable manifold and for each $y\in R$, there exists a unique $\gamma^{u}_{y}\in\Gamma^{u}$ such that $y\in\gamma^{u}_{y}$
  • for every $\gamma^{s}\in\Gamma^{s}$ and $\gamma^{u}\in\Gamma^{u}$ one has $(\gamma^{s}\cap\gamma^{u})=1$.
Given a rectangle $R$ and $x,y\in R$, define the Smale bracket
$x,y]_{\Theta}:=\gamma^{s}_{x}\cap\gamma^{u}_{y},$
which is a single point of $R$ by the last item. We also define the sets
$W^{s}(x,R):=\gamma^{s}_{x}\cap R\ \text{and }W^{u}(x,R):=\gamma^{u}_{x}\cap R.$
Figure 1. Illustration of a rectangle $R$ (shown as green dots) as the intersection of segments of stable and unstable manifolds (blue and red curves, respectively).
@@ 3. Topological Markov shifts
equipped with the metric $d(v,u)=\{\exp(-\min\{|n|;v_{n}\neq u_{n}\})$, and $\sigma:\Sigma\to\Sigma$ is the left shift map $\sigma((v_{i})_{i\in\mathbb{Z}})=(v_{i+1})_{i\in\mathbb{Z}}$. Given a topological Markov shift $\Sigma$ associated to a graph $\mathscr{G}=(\mathscr{V},\mathscr{E})$ we define the set
@@
In the following, we recall the definition of a Markov partition.
Definition 3.2 (Markov partition).
Let $\theta:\Omega\to\Omega$ be a $\mathcal{C}^{1+\alpha}$ dynamical system. A finite family $\mathscr{P}=\{P_{1},\dots,P_{k}\}$ of rectangles with local product structure (see Definition 2.11) is a Markov partition if:
@@
    for each $i\in\{1,\ldots,k\}$
    $P_{i}=\overline{\operatorname{int}(P_{i})}\ \text{and}\ \operatorname{int}(P_{i})\cap\operatorname{int}P_{j}=\varnothing$
    if $i\neq j$, where the interior and closure are being taken in the induced topology of $NW(\theta)$, where $NW(\theta)$ denotes the non-wandering set of $\theta:\Omega\to\Omega$;
@@
    whenever $\omega\in\operatorname{int}P_{i}$ and $\theta(\omega)\in\operatorname{int}P_{j}$,
    $\theta\big(W^{s}(\omega,P_{i})\big)\subset W^{s}(\theta(\omega),P_{j})\text{ and }\theta\big(W^{u}(\omega,P_{i})\big)\supset W^{u}(\theta(\omega),P_{j}).$
@@
    define the directed graph $\mathscr{G}_{b}=(\mathscr{V}_{b},\mathscr{E}_{b})$ with $\mathscr{V}_{b}=\mathscr{P}$ and edges
    $\mathscr{E}_{b}=\big\{P_{i}\to P_{j};\ \operatorname{int}P_{i}\cap\theta^{-1}(\operatorname{int}P_{j})\neq\varnothing\,\big\},$
    and let $\Sigma_{b}=\Sigma_{b}(\mathscr{G}_{b})$ be the associated subshift of finite type. The coding map
    $\pi_{b}:(P_{i_{n}})_{n\in\mathbb{Z}}\ \mapsto\ \bigcap_{n\in\mathbb{Z}}\theta^{-n}\big(P_{i_{n}}\big)$ (3)
    is well defined, Hölder continuous, and satisfies the semiconjugacy $\pi_{b}\circ\sigma=\theta\circ\pi_{b}$.
  • (4)
    the map $\pi_{b}$ is surjective onto $NW(\theta)$ and there exists $k_{\mathscr{P}}\in\mathbb{N}$ such that $\pi_{b}^{-1}(\omega)\leq k_{\mathscr{P}}$ for every $\omega\in NW(\theta)$.
Theorem 3.3 ([Bowen70], [ParryPollicott1990, Appendix III]).
Every $\mathcal{C}^{1+\alpha}$ Axiom A diffeomorphism $\theta:\Omega\to\Omega$ admits a finite Markov partition; in particular, so does every Anosov diffeomorphism.
Definition 3.4.
Let $\theta:\Omega\to\Omega$ be an Anosov diffeomorphism and let $\mathbb{P}$ be $\theta$-invariant ergodic probability measure on $\Omega$. Given a Markov partition $\mathscr{P}=\{P_{1},\ldots,P_{\kappa}\}$ for $\theta$ we denote
$\partial\mathscr{P}:=\bigcup_{P\in\mathscr{P}}\partial P.$
It is convenient to decompose the boundary into its stable and unstable parts:
$\partial^{s}\mathscr{P}:=\bigcup_{P\in\mathscr{P}}\partial^{s}P,\ \partial^{u}\mathscr{P}:=\bigcup_{P\in\mathscr{P}}\partial^{u}P,$
where for each $P\in\mathscr{P}$,
  • $\partial^{s}P:=\{x\in P:x\notin\operatorname{int}_{W^{u}(x)}(W^{u}(x,P))\}$; and
  • $\partial^{u}P:=\{x\in P:x\notin\operatorname{int}_{W^{s}(x)}(W^{s}(x,P))\},$
where interiors taken in the relative topologies of $W^{s}_{\mathrm{loc}}(x)$ and $W^{u}_{\mathrm{loc}}(x)$, respectively (see [BowenBook, Lemma 3.11]). Observe that $\mathbb{P}$ is ergodic, we have that $\mathbb{P}(\partial\mathscr{P})\in\{0,1\}$ (see [BowenBook, Proposition 3.5]).
The next statement, taken from [buzzi2025strongpositiverecurrenceexponential, Theorem 10.5 and Section 10], is a higher-dimensional generalisation of Sarig’s celebrated theorem [Sarig2013, Theorem 1.3] (see also [BenOvadia2018, Theorem 0.1]). We mention that the original result presented in [buzzi2025strongpositiverecurrenceexponential] is more general; however, the presented version is enough for our purposes.
Theorem 3.5 ([buzzi2025strongpositiverecurrenceexponential, Theorem 10.5 and Section 10], see also [Sarig2013, Proposition 11.5] and [BenOvadia2018, §6]).
Let $\Theta:E\to E$ be a $\mathcal{C}^{1+\alpha}$ diffeomorphism on a compact smooth manifold $E$. Let $\mathcal{X}$ be either a Borel homoclinic class $H$ or $E$. Then, for each $\chi>0$, there exists a countable collection of sets $\mathscr{R}=\{R_{i}\}_{i\in I}$, which we call a $\chi$-Sarig partition, with the following properties:
  • Each $R_{i}$ is a rectangle with local product structure (see Definition 2.11);
  • whenever $x\in R_{i}$ and $\Theta(x)\in R_{j}$, then
    $\Theta\big(W^{s}(x,R_{i})\big)\subset W^{s}(\Theta(x),R_{j})\ \text{and}\ \Theta\big(W^{u}(x,R_{i})\big)\supset W^{u}(\Theta(x),R_{j}).$
Let the topological Markov shift $\Sigma=\Sigma(\mathscr{G})$ be induced by the directed graph $\mathscr{G}=(\mathscr{V},\mathscr{E})$ with $\mathscr{V}:=\mathscr{R}$ and
$\mathscr{E}=\big\{\,R_{i}\to R_{j};\ R_{i}\cap\Theta^{-1}(R_{j})\neq\varnothing\,\big\}.$
Then:
@@
    $\pi:\ (R_{i})_{i\in\mathbb{Z}}\in\Sigma\ \mapsto\ \bigcap_{n\in\mathbb{N}}\,\overline{\bigcap_{-n\leq i\leq n}\Theta^{-i}(R_{i})}\in\mathcal{X}$
    is well-defined and Hölder continuous. Moreover, $\Theta\circ\pi=\pi\circ\sigma$ and $\pi:\Sigma^{\#}\to\mathcal{X}$ is finite-to-one (see Definition 3.1).
@@
    $\Sigma$ is a locally compact countable-state Markov shift and $\mu(\pi(\Sigma))=\mu(\pi(\Sigma^{\#}))=1$. If $\mathcal{X}$ is a Borel homoclinic class, $\Sigma$ can be chosen irreducible.
@@
    Every ergodic $\sigma$-invariant probability measure $\widehat{\mu}$ on $\Sigma$ projects to an ergodic $\Theta$-invariant measure $\mu:=\pi_{*}\widehat{\mu}$ with $\mu(\mathcal{X})=1$ and $h_{\mu}(\Theta)=h_{\widehat{\mu}}(\sigma)$.
@@
    For every $\chi$-hyperbolic $\Theta$-invariant probability measure $\mu$ on $E$ there exists a $\sigma$-invariant probability measure $\widehat{\mu}$ on $\Sigma$ such that $\pi_{*}\widehat{\mu}=\mu$.
We now state the main technical result on which our analysis is based. It refines Theorem 3.5 in the setting of skew products.
@@
    $\Theta$ is a skew product i.e. $\Theta(\omega,x)=(\theta(\omega),T_{\omega}(x))$;
@@
    $\theta:\Omega\to\Omega$ is an Anosov map and $\mathbb{P}$ a $\theta$-invariant ergodic probability measure with full support.
Let $\mathcal{X}\subset\Omega\times M$ be equal to a Borel homoclinic class $H$ or $\Omega\times M$. Then, for every $\chi>0$ there exists a locally compact topological Markov shift $\Sigma_{s}$, a subshift of finite type $\Sigma_{b}$ and Hölder continuous maps $\pi_{s}:\Sigma_{s}\to\mathcal{X}$ and $\pi_{b}:\Sigma_{b}\to\Omega$ such that
@@
    for every $\chi$-hyperbolic $\Theta$-invariant probability $\mu$ such that $\mu(\mathcal{X})=1$ with $(\mathrm{proj}_{\Omega})_{*}\mu=\mathbb{P}$, one has $\mu\big(\pi_{s}[\Sigma_{s}]\big)=\mu\big(\pi_{s}[\Sigma_{s}^{\#}]\big)=1$};
@@
    the maps $\pi_{s}\big|_{\Sigma_{s}^{\#}}:\Sigma_{s}^{\#}\to\mathcal{X}$ and $\pi_{b}:\Sigma_{b}\to\Omega$ are finite-to-one;
@@
    Moreover, $\widehat{\mathbb{P}}\left[\mathrm{proj}_{\Sigma_{b}}(\Sigma_{s})\right]=1$ where $\widehat{\mathbb{P}}$ is the unique measure in $\mathcal{M}_{e}(\Sigma_{b})$ such that $(\pi_{b})_{*}\widehat{\mathbb{P}}$$\mathbb{P}$.
@@
    if $\mathcal{X}$ is a Borel homoclinic class and there exists a $\chi$-hyperbolic $\Theta$-invariant ergodic probability measure $\mu$ with $\operatorname{supp}\mu\subset\mathcal{X}$ and $(\mathrm{proj}_{\Omega})_{*}\mu=\mathbb{P}$, then the coding $(\Sigma_{s},\sigma_{s})$ can be chosen to be an irreducible topological Markov shift.
@@
$\mathcal{X}\subset\Omega\times M$$\mathcal{X}\subset\Omega\times M$$\Sigma$$\Sigma$$\Sigma_{s}$$\Sigma_{s}$$\Sigma_{b}$$\Sigma_{b}$$\Omega$$\Omega$$\scriptstyle{{\rm proj}_{\Omega}}$$\scriptstyle{\Theta}$$\scriptstyle{{\rm proj}_{\Omega}}$$\scriptstyle{\sigma}$$\scriptstyle{\pi}$$\scriptstyle{\pi}$$\scriptstyle{\pi_{s}}$$\scriptstyle{{\rm proj}_{\Sigma_{b}}}$$\scriptstyle{\sigma_{s}=\sigma_{b}\times\sigma}$$\scriptstyle{{\rm proj}_{\Sigma}}$$\scriptstyle{{\rm proj}_{\Sigma}}$$\scriptstyle{\pi_{s}}$$\scriptstyle{{\rm proj}_{\Sigma_{b}}}$$\scriptstyle{\sigma_{b}}$$\scriptstyle{\pi_{b}}$$\scriptstyle{\pi_{b}}$$\scriptstyle{\theta}$
@@
Let $\chi>0$. Applying Theorem 3.5 to the skew product $\Theta:\Omega\times M\to\Omega\times M$ and $\mathcal{X}$ be either equal to Borel homoclinic class $H$ or $\Omega\times M$, we obtain a topological Markov shift $(\Sigma,\sigma)$ with the properties asserted therein, and let $\mathscr{R}$ be its $\chi$-Sarig partition. We assume $(\Sigma,\sigma)$ to be irreducible if $\mathcal{X}$ is a Borel homoclinic class. Since $\theta:\Omega\to\Omega$ is a transitive Anosov diffeomorphism and $\mathbb{P}$ is a measure with full support, there exists a subshift of finite type $(\Sigma_{b},\sigma_{b})$ associated with a Markov partition $\mathscr{P}$ as in Definition 3.2. Since $\mathrm{supp}\,\mathbb{P}=\Omega$ we have that $\mathbb{P}(\partial\mathscr{P})=0$ (see Definition 3.4). Therefore
$\mathbb{P}\left[\omega\in\Omega;\,\theta^{i}(\omega)\in\partial\mathscr{P}\ \text{for some }i\in\mathbb{Z}\right]$ $=\mathbb{P}\left[\bigcup_{i\in\mathbb{Z}}\theta^{-i}(\partial\mathscr{P})\right]$
$\leq\sum_{i\in\mathbb{Z}}\mathbb{P}\left[\theta^{-i}(\partial\mathscr{P})\right]=0.$ (4)
Let $\pi_{b}$ be as in (3) and recall that
$k_{\mathscr{P}}:=\sup_{\omega\in\Omega}\#\pi_{b}^{-1}(\omega)<\infty.$
Let $\mathscr{G}_{s}:=(\mathscr{V}_{s},\mathscr{E}_{s})$ be the graph with vertices
$\mathscr{V}_{s}:=\{(P,R);\,R\in\mathscr{R},P\in\mathscr{P}\ \text{and }R\cap(\operatorname{int}P\times M)\neq\varnothing\},$
and edges
$\mathscr{E}_{s}=\left\{(P,R)\to(P^{\prime},R^{\prime})\;\middle|\;\begin{aligned} &(P,R),(P^{\prime},R^{\prime})\in\mathscr{V}_{s},\ \text{and }\\ &\Theta^{-1}[R^{\prime}\cap(\operatorname{int}P^{\prime}\times M)]\cap R\cap(\operatorname{int}P\times M)\neq\varnothing\end{aligned}\right\}.$
Define $(\Sigma_{s},\sigma_{s}):=(\Sigma_{s}(\mathscr{G}_{s}),\sigma_{s})$. Observe that for each $(P,R)\in\mathscr{V}_{s}$
$\{(P^{\prime},R^{\prime})\in\mathscr{V}_{s};(P,R)\to(P^{\prime},R^{\prime})\in\mathscr{E}_{s}\}\leq\#\{R^{\prime}\in\mathscr{V}_{s};R\to R^{\prime}\in\mathscr{E}\}\cdot\#\mathscr{P}<\infty,$